OpenStudy (anonymous):

Simplify the expression √-1/(4-3i)(2+5i)

OpenStudy (anonymous):

\[\sqrt{(-1)/(8+14i-15i^2)}\] =\[\sqrt{(-1)/(8+14i+15)}\] =\[\sqrt{(-1)/(23+14i)}\]

OpenStudy (anonymous):

Um...keep going

OpenStudy (anonymous):